High Desert Potteryworks makes a variety of pottery products that it sells to retailers. The company uses a job-order costing system in which departmental predetermined overhead rates are used to apply manufacturing overhead cost to jobs. The predetermined overhead rate in the Molding Department is based on machine-hours, and the rate in the Painting Department is based on direct labor- hours. At the beginning of the year, the compány provided the following estimates: Department Molding Painting 12,00 60,80e 70,000 8,0e0 Direct labor-hours Machine-hours Fixed manufacturing overhead cost 497,000 $615,88 Variable manufacturing overhead per machine-hour Variable manufacturing overhead per direct labor-hour $ 1.50 Job 205 was started on August 1 and completed on August 10. The companys cost records show the following information concerning the job: Departmen Molding Painting Direct 1abor-hours Machine-hours Direct materials Direct labor cost 30 116 84 20 $770 $1,332 $525 $1,470 Required: 1. Compute the predetermined oyerhead rates used in the Molding Department and the Painting Department. 2. Compute the total overhead cost applied to Job 205. 3-a. What would be the total manufacturing cost recorded for Job 205? 3-b. If the job contained 50 units, what would be the unit product cost?
Required: 1. Compute the predetermined overhead rates used in the Molding Department and the Painting Department. 2. Compute the total overhead cost applied to Job 205. 3-a. What would be the total manufacturing cost recorded for Job 205? 3-b. If the job contained 50 units, what would be the unit product cost? Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3A Required 38 Compute the predetermined overhead rates used in the Molding Department and the Painting Department. (Round your answers to 2 decimal places) Predetermined Overhead Rate Molding Department Painting Department per MH per DLH Required 1 Required 2>
Complete this question by entering your answers in the tabs below Required 1Required 2 Required 3A Required 3B Compute the total overthead cost applied to Job 205. (Round your intermediate calculations to 2 decilmal places,) Total overhead cost Required 1 Required 3A
media%2F8a5%2F8a545b9a-a573-4d17-bb1f-26
media%2F07f%2F07fd05c9-97c6-412d-a235-ff
Answer #1

a) Predetermine overhead rate :

Molding department = (497000/70000)+1.50 = $8.60 per mh

Painting department = (615000/60000)+2.00 = $12.25 per dlh

b) Total overhead cost = (110*8.60+84*12.25) = $1975

c) Calculate total product cost

Molding Painting Total
Direct material 770 1332 2102
Direct labor 525 1470 1995
Overhead applied 946 1029 1975
Total 2241 3831 6072

d) Unit product cost = 6072/50 = $121.44 per unit